# Stellar Mass Segregation in Dark Matter Halos

### v1

@@ 4.4 The effect of tides
The simulation results discussed in Sec. 4.1 assume a static dark matter potential surrounding the stellar populations. For the faint stellar systems UMa3/U1, Del1 this assumption is unlikely to hold: located at galactocentric distances similar to that of the sun, these systems will be subject to the tidal field of the Milky Way. For the case of UMa3/U1 on an orbit with a pericentre of $r_{\mathrm{peri}}=13\,\mathrm{kpc}$, a dark matter halo hosting UMa3/U1 could have been tidally stripped to $1/10^{4}$ of its original mass (see Fig. 7 in Errani et al. 2024b).
To model the effect of tides, in the following, we will slowly decrease the mass of the surrounding dark matter halo and adjust its scale radius according to the empirical tidal evolutionary tracks for Hernquist models (see Errani et al. 2018 Table. A1 and Fig. A1 for details). Specifically, we model the evolution of the subhalo total mass as $M_{\mathrm{sub}}/M_{\mathrm{sub0}}=\exp(-t/\tau)$ where $\tau$ is chosen so that over $10\,\mathrm{Gyr}$ of evolution, $M_{\mathrm{sub}}/M_{\mathrm{sub0}}$ decreases to $1/10^{4}$. The scale radius is adjusted through $r_{\mathrm{sub}}/r_{\mathrm{sub0}}\approx(M_{\mathrm{sub}}/M_{\mathrm{sub0}})^{\beta}$ with $\beta=0.48$.
Taking UMa3/U1 as an example, figure 6 shows the results of this experiment. A red curve shows the evolution of the (median) half-light radius $r_{\mathrm{h}}$ and velocity dispersion $\langle v^{2}\rangle^{1/2}$ of the population of low-mass stars, normalized to the respective initial values. A grey curve shows the equivalent evolution for the population of high-mass stars. As previously discussed, the stellar populations mass segregate. Here, in addition, the lowering of the background dark matter potential drives an adiabatic expansion of the stellar components along curves of $r_{\mathrm{h}}\langle v^{2}\rangle^{1/2}=\text{const}$ (black dashed curves, see Errani et al. 2025 for detailed discussion of this adiabatic expansion). Note that, in the size–velocity dispersion plane, the dynamical effects of tides and mass segregation are orthogonal: mass segregation results in an overall expansion and heating (contraction and cooling) of the population of low-mass (high-mass) stars, whereas tides drive an adiabatic expansion and cooling of both components. For the case of the population of high-mass stars, these two processes compete in driving the half-light radius of the stellar population. Crucially, after $10\,\mathrm{Gyr}$ of evolution, the ratio of the half-light radii of the low-mass and high-mass populations are virtually identical between the models which include adiabatic expansion through tides (red and grey lines), and the model run in isolation (blue and dark grey lines). For the case of tidal fields that cause an adiabatic expansion of the stellar component within the power-law cusp of the underlying halo, the ratio of half-light radii shown in Fig. 5 will hence hold independently of whether the system has experienced tides, or not. This finding is consistent with the analytical estimate of the relaxation time in Eq. 6: for constant $N_{\star}$ and $m_{\star}$, the relaxation time scales as $T_{\mathrm{rel}}\propto(M_{\mathrm{h}}r_{\mathrm{h}})^{3/2}$. The product of enclosed mass and half-light radius is approximately conserved during adiabatic expansion, $M_{\mathrm{h}}r_{\mathrm{h}}\propto\langle v^{2}\rangle^{1/2}r_{\mathrm{h}}\approx\text{const}$ (see e.g. Errani et al. 2025 for details). Hence, the relaxation time (Eq. 6) and the binary fraction (Eq. 9) remains virtually unaffected by the adiabatic expansion of the stellar components.
To illustrate the potential landscape and to provide further intuition for the expected evolution in the size–velocity dispersion plane, black curves show the velocity dispersion of a tracer population subject to the combined potential of dark matter and stars, as computed [^8] Velocity dispersions are additive, and we can compute separately the dispersions expected for a stellar component due to (a) the potential of another stellar component, (b) its self gravity, and (c) the Hernquist dark matter halo. In each case, we will use Eq. 3 of Errani et al. (2018) for the calculation of the velocity dispersion. The velocity dispersion of a (mass-less) exponential stellar tracer (Eq. 7) with scale radius $r_{i}$ in the potential of another stellar component of mass $M_{j}$ and scale radius $r_{j}$ reads $\langle v^{2}_{\text{a}}\rangle=GM_{j}/(2r_{i})\left(1+4r_{j}/r_{i}\right)\left(1+r_{j}/r_{i}\right)^{-4}$ (10) which for the self-gravitating case, $r_{i}=r_{j}$ and $M_{i}=M_{j}$ reduces to $\langle v^{2}_{\text{b}}\rangle=(5/96)~{}GM_{i}/r_{i}$. For an exponential stellar tracer with scale radius $r_{i}$ embedded in a Hernquist potential (Eq. 8), writing for short $x=r_{\mathrm{sub}}/r_{i}$, we find $\langle v^{2}_{\text{c}}\rangle=\frac{GM_{\mathrm{sub}}}{2r_{i}}\left[2\!-\!(1\!+\!x)^{2}\!-\!x^{2}(x\!+\!3)\exp(x)\mathrm{Ei}(-x)\right]$ (11) which for deeply embedded systems, $x\gg 1$, approaches the power-law $\langle v^{2}_{\mathrm{c}}\rangle\approx 3GM_{\mathrm{sub}}r_{i}/r_{\mathrm{sub}}^{2}$. Note that for large values of $x$, numerical evaluations of Eq. 11 may be facilitated by expressing the product $\exp(x)\mathrm{Ei}(-x)$ by its approximation as a Laurent series $-x^{-1}+x^{-2}-2x^{-3}+6x^{-4}-24x^{-5}+\mathcal{O}(x^{-6})$. from the Virial theorem (see e.g. Amorisco & Evans 2012, Errani et al. 2018). This simple calculation accurately predicts the velocity dispersion of the mass-segregated populations.
The model for tides employed here does not account for any stellar mass loss, but merely models the response of the stars to the evolving background potential. This is a modelling choice motivated by the fact that asymptotic remnant state (Errani & Navarro, 2021) of a cold dark matter halo on the orbit of UMa3/U1 has a tidal radius that is substantially larger than the half-light radius of UMa3/U1. For stronger tidal fields that result in the tidal stripping of stars, this assumption will not hold. As mass segregation drives low-mass stars to less-bound orbits, and high-mass stars to more bound ones, in a mass segregated system, tides would first strip the population of low-mass stars. This cloud result in the existence of a population of dark matter subhalos that hosts high-mass stars or their remnants at their centers: black holes surrounded by dark matter subhalos. Their mergers would in turn facilitate the formation of massive black holes in the centres of dark matter-dominated systems. This will be explored in future contributions.
@@ 5 Conclusions
Summary. In the present work, we show that effects of collisional relaxation may play a substantial role for the dynamical evolution of a stellar component even in dark matter-dominated system. This is of particular relevance for old stellar systems with short crossing times, where small collisional perturbations can accumulate over the course of several gigayears. We show that for such systems, collisional relaxation drives stellar mass segregation and the dynamical formation of binaries even in presence of a gravitationally dominant smooth dark matter component. Our results hence call for caution when using stellar mass segregation as a litmus test for the absence of dark matter in ambiguous stellar clusters (Kim et al., 2015; Baumgardt et al., 2022; Simon et al., 2024; Zaremba et al., 2025) and tidal streams (Errani et al., 2022). Detailed modelling of the relaxation time scale and the Poisson noise floor is required to put constraints on the dark matter content of a stellar system through the observable signatures of mass segregation.
@@
Caveats. Our models make various simplifying assumptions. Most crucially, the stellar populations are modelled as a two-component system of high- and low mass stars, both initially sharing the same half-light radius. The relaxation time of a stellar system sensitively depends on its stellar mass function and the abundance of high-mass stars. Our models do not include high-mass stellar remnants which may constitute a source of additional collisional perturbations which could amplify and speed up mass segregation. In that regard we believe our modelling choices to be conservative, putting a lower bound on the amount of mass segregation that is to be expected in presence of a smooth dark matter subhalo.
@@
Furthermore, our models are tailored to describe systems that remain dark matter-dominated throughout their evolution, and neglect the dynamical effects of the stars on the dark matter distribution. The same fluctuating tidal field which drives stellar mass segregation is likely to affect the dark matter as well, particularly in systems that are initially or become baryon-dominated. A detailed study of this effect is computationally expensive and beyond the scope of the current paper.
@@
Outlook. The models developed in this work are motivated by the recent discovery of ambiguous stellar systems at the interface of the dwarf galaxy- and globular cluster regimes, nevertheless they can also find application in understanding the dynamical processes in the central regions of dwarf spheroidal- and ultra faint galaxies. While their half-mass relaxation times may exceed the age of the universe, mass segregation plausibly still plays a role in their centres, where dynamical time scales and stellar densities are similar those of the systems studied here. In addition to the effects of dynamical friction by the dark matter component, stellar mass segregation may further enhance the clustering of massive stars and their remnants in the centres of dwarf galaxies, plausibly playing a role in setting their merger rates and a potential accompanying gravitational wave signal. This in turn would facilitate the formation of massive black holes in the centres of dark matter-dominated dwarf galaxies. A quantitative analysis of this effect requires a detailed modelling of the stellar mass function beyond the two-component setup used in the present work, as well as a detailed modelling of the dynamical response of the dark matter cusp, which we defer to subsequent study.
@@ Acknowledgements
The authors would like to thank Anna Lisa Varri, Rodrigo Ibata, Giacomo Monari and Josh Simon for stimulating discussions. RE and MW acknowledge support from the National Science Foundation (NSF) grant AST-2206046. Support for program JWST-AR-02352.001-A was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This material is based upon work supported by the National Aeronautics and Space Administration under Grant/Agreement No. 80NSSC24K0084 as part of the Roman Large Wide Field Science program funded through ROSES call NNH22ZDA001N-ROMAN.

### v2

@@ 4.4 The effect of tides
The simulation results discussed in Sec. 4.1 assume a static dark matter potential surrounding the stellar populations. For the faint stellar systems UMa3/U1, Del1 this assumption is unlikely to hold: located at galactocentric distances similar to that of the Sun, these systems will be subject to the tidal field of the Milky Way. For the case of UMa3/U1 on an orbit with a pericentre of $r_{\mathrm{peri}}=13\,\mathrm{kpc}$, a dark matter halo hosting UMa3/U1 could have been tidally stripped to $1/10^{4}$ of its original mass (see Fig. 7 of Errani et al. 2024b).
@@
To model the effect of tides, in the following, we will slowly decrease the mass of the surrounding dark matter halo and adjust its scale radius according to the empirical tidal evolutionary tracks for Hernquist models (see Errani et al. 2018 Table A1 and Fig. A1 for details). Specifically, we model the evolution of the subhalo total mass as $M_{\mathrm{sub}}/M_{\mathrm{sub0}}=\exp(-t/\tau)$ where $\tau$ is chosen so that over $10\,\mathrm{Gyr}$ of evolution, $M_{\mathrm{sub}}/M_{\mathrm{sub0}}$ decreases to $1/10^{4}$. The scale radius is adjusted through $r_{\mathrm{sub}}/r_{\mathrm{sub0}}\approx(M_{\mathrm{sub}}/M_{\mathrm{sub0}})^{\beta}$ with $\beta=0.48$.
Taking UMa3/U1 as an example, Fig. 6 shows the results of this experiment. A red curve shows the evolution of the (median) half-light radius $r_{\mathrm{h}}$ and velocity dispersion $\langle v^{2}\rangle^{1/2}$ of the population of low-mass stars, normalized to the respective initial values. A grey curve shows the equivalent evolution for the population of high-mass stars. As previously discussed, the stellar populations segregate in mass. Here, in addition, the lowering of the background dark matter potential drives an adiabatic expansion of the stellar components along curves of $r_{\mathrm{h}}\langle v^{2}\rangle^{1/2}=\text{const}$ (black dashed curves, see Errani et al. 2025 for detailed discussion of this adiabatic expansion). Note that, in the size–velocity dispersion plane, the dynamical effects of tides and mass segregation are orthogonal: mass segregation results in an overall expansion and heating (contraction and cooling) of the population of low-mass (high-mass) stars, whereas tides drive an adiabatic expansion and cooling of both components. For the case of the population of high-mass stars, these two processes compete in driving the half-light radius of the stellar population. Crucially, after $10\,\mathrm{Gyr}$ of evolution, the ratio between the half-light radii of the low-mass and high-mass populations is virtually identical in the models that include adiabatic expansion through tides (red and grey lines), and in the models run in isolation (blue and dark grey lines). For the case of tidal fields that cause an adiabatic expansion of the stellar component within the power-law cusp of the underlying halo, the ratio of half-light radii shown in Fig. 5 will hence hold independently of whether the system has experienced tides or not. This finding is consistent with the analytical estimate of the relaxation time in Eq. 6: for constant $N_{\star}$ and $m_{\star}$, the relaxation timescales as $T_{\mathrm{rel}}\propto(M_{\mathrm{h}}r_{\mathrm{h}})^{3/2}$. The product of enclosed mass and half-light radius is approximately conserved during adiabatic expansion, $M_{\mathrm{h}}r_{\mathrm{h}}\propto\langle v^{2}\rangle^{1/2}r_{\mathrm{h}}\approx\text{const}$ (see, e.g., Errani et al. 2025 for details). Hence, the relaxation time (Eq. 6) and the binary fraction (Eq. 9) remain virtually unaffected by the adiabatic expansion of the stellar components.
To illustrate the potential landscape and to provide further intuition for the expected evolution in the size–velocity dispersion plane, black curves show the velocity dispersion of a tracer population subject to the combined potential of dark matter and stars, as computed [^10] Velocity dispersions are additive, and we can compute separately the dispersions expected for a stellar component due to (a) the potential of another stellar component, (b) its self gravity, and (c) the Hernquist dark matter halo. In each case, we will use Eq. 3 of Errani et al. (2018) for the calculation of the velocity dispersion. The velocity dispersion of a (massless) exponential stellar tracer (Eq. 7) with scale radius $r_{i}$ in the potential of another stellar component of mass $M_{j}$ and scale radius $r_{j}$ reads $\langle v^{2}_{\text{a}}\rangle=\frac{GM_{j}}{2r_{i}}\left(1+4r_{j}/r_{i}\right)\left(1+r_{j}/r_{i}\right)^{-4}$ (10) which for the self-gravitating case, $r_{i}=r_{j}$ and $M_{i}=M_{j}$ reduces to $\langle v^{2}_{\text{b}}\rangle=(5/96)~GM_{i}/r_{i}$. For an exponential stellar tracer with scale radius $r_{i}$ embedded in a Hernquist potential (Eq. 8), writing for short $x=r_{\mathrm{sub}}/r_{i}$, we find $\langle v^{2}_{\text{c}}\rangle=\frac{GM_{\mathrm{sub}}}{2r_{i}}\left[2\!-\!(1\!+\!x)^{2}\!-\!x^{2}(x\!+\!3)\exp(x)\mathrm{Ei}(-x)\right]$ (11) which for deeply embedded systems, $x\gg 1$, approaches the power law $\langle v^{2}_{\mathrm{c}}\rangle\approx 3GM_{\mathrm{sub}}r_{i}/r_{\mathrm{sub}}^{2}$. Note that for large values of $x$, numerical evaluations of Eq. 11 may be facilitated by expressing the product $\exp(x)\mathrm{Ei}(-x)$ by its approximation as a Laurent series: $x^{-1}+x^{-2}-2x^{-3}+6x^{-4}-24x^{-5}+\mathcal{O}(x^{-6})$. from the virial theorem (see, e.g., Amorisco & Evans 2012; Errani et al. 2018). This simple calculation accurately predicts the velocity dispersion of the mass-segregated populations.
The model for tides employed here does not account for any stellar mass loss, but merely models the response of the stars to the evolving background potential. This is a modeling choice motivated by the fact that the asymptotic remnant state (Errani & Navarro, 2021) of a cold dark matter halo on the orbit of UMa3/U1 has a tidal radius that is substantially larger than the half-light radius of UMa3/U1. For stronger tidal fields that result in the tidal stripping of stars, this assumption will not hold. As mass segregation drives low-mass stars to less bound orbits, and high-mass stars to more bound ones, in a mass-segregated system, tides would first strip the population of low-mass stars. This could result in the existence of a population of dark matter subhalos that hosts high-mass stars or their remnants at their centers: black holes surrounded by dark matter subhalos. Their mergers would in turn facilitate the formation of massive black holes in the centres of dark matter-dominated systems. This will be explored in future contributions.
@@ 5 Conclusions
Summary. In the present work, we show that effects of collisional relaxation may play a substantial role in the dynamical evolution of a stellar component even in a dark matter-dominated system. This is of particular relevance for old stellar systems with short crossing times, where small collisional perturbations can accumulate over the course of several gigayears. We show that for such systems, collisional relaxation drives stellar mass segregation and the dynamical formation of binaries even in the presence of a gravitationally dominant smooth dark matter component. Our results hence call for caution when using stellar mass segregation as a litmus test for the absence of dark matter in ambiguous stellar clusters (Kim et al., 2015; Baumgardt et al., 2022; Simon et al., 2024; Zaremba et al., 2025) and tidal streams (Errani et al., 2022). Detailed modeling of the relaxation timescale and the Poisson noise floor is required to put constraints on the dark matter content of a stellar system through the observable signatures of mass segregation.
@@
Caveats. Our models make various simplifying assumptions. Most crucially, the stellar populations are modeled as a two-component system of high- and low-mass stars, the two initially sharing the same half-light radius. The relaxation time of a stellar system depends sensitively on its stellar mass function and the abundance of high-mass stars. Our models do not include high-mass stellar remnants, which may constitute a source of additional collisional perturbations that could amplify and speed up mass segregation. In that regard we believe our modeling choices to be conservative, putting a lower bound on the amount of mass segregation that is to be expected in the presence of a smooth dark matter subhalo.
@@
Furthermore, our models are tailored to describe systems that remain dark matter-dominated throughout their evolution, and neglect the dynamical effects of the stars on the dark matter distribution. The same fluctuating tidal field that drives stellar mass segregation is likely to affect the dark matter as well, particularly in systems that are initially or become baryon-dominated. A detailed study of this effect is computationally expensive and beyond the scope of the current paper.
@@
Outlook. The models developed in this work are motivated by the recent discovery of ambiguous stellar systems at the interface of the dwarf galaxy and globular cluster regimes; nevertheless they can also find application in understanding the dynamical processes in the central regions of dwarf spheroidal and ultrafaint galaxies. While their half-mass relaxation times may exceed the age of the Universe, mass segregation plausibly still plays a role in their centres, where dynamical timescales and stellar densities are similar to those of the systems studied here. In addition to the effects of dynamical friction by the dark matter component, stellar mass segregation may further enhance the clustering of massive stars and their remnants in the centres of dwarf galaxies, plausibly playing a role in setting their merger rates and a potential accompanying gravitational wave signal. This in turn would facilitate the formation of massive black holes in the centres of dark matter-dominated dwarf galaxies (see, e.g., Bustamante-Rosell et al. 2021 for the case of Leo I). A quantitative analysis of this effect requires a detailed modeling of the stellar mass function beyond the two-component setup used in the present work, as well as a detailed modeling of the dynamical response of the dark matter cusp, which we defer to subsequent study.
@@ Acknowledgements
The authors would like to thank Anna Lisa Varri, Rodrigo Ibata, Giacomo Monari, and Josh Simon for stimulating discussions. RE and MW acknowledge support from the National Science Foundation (NSF) grant AST-2206046. Support for program JWST-AR-02352.001-A was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This material is based upon work supported by the National Aeronautics and Space Administration under Grant/Agreement No. 80NSSC24K0084 as part of the Roman Large Wide Field Science program funded through ROSES call NNH22ZDA001N-ROMAN.
Appendix A Supplementary animated figures
This appendix contains brief descriptions of the animated figures 7 and 8, available in the arXiv ancillary files.
Figure 7: Animated figure supplementing the top panel of Fig. 2. The video shows the time evolution of a population of high- and low-mass stars (black and blue points, respectively) in the presence of a smooth dark matter potential. Circles indicate the median half-light radii measured from a sample of $N$-body realizations (see Sec. 4.1). File size $4.9\,\mathrm{MB}$, dimensions $1920\times 1920\,\mathrm{px}^{2}$, duration $40\,\mathrm{s}$. The animation shows the evolution from 0 to $10\,\mathrm{Gyr}$. The playback speed is increased between seconds $14$ ($0.5\,\mathrm{Gyr}$) and $26$ ($9.5\,\mathrm{Gyr}$) for the sake of brevity. The animated figure is available in the arXiv ancillary files.
Figure 8: Like Fig. 7, but highlighting the dynamical formation and subsequent disruption of a stellar binary (red circles). File size $1.8\,\mathrm{MB}$, dimensions $1920\times 1920\,\mathrm{px}^{2}$, duration $28\,\mathrm{s}$. The animation shows the evolution from $7.75$ to $8.25\,\mathrm{Gyr}$. The animated figure is available in the arXiv ancillary files.